OpenStudy (anonymous):

how do you solve x=x^2+1/2

OpenStudy (kittiwitti1):

\[x = x ^{2} + \frac{ 1 }{ 2 }\]Move the x to the right side.\[0 = x ^{2} + \frac{ 1 }{ 2 } - x\]\[x ^{2} - x + \frac{ 1 }{ 2 } = 0\] *** Use the quadratic equation to solve.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

she is right :) i agree with her answer.